# DiSCo Meets LLMs: A Unified Approach for Sparse Retrieval and Contextual Distillation in Conversational Search

### v1

@@ Acknowledgements.
- Lassance et al. (2024a) Carlos Lassance, Hervé Dejean, Stéphane Clinchant, and Nicola Tonellotto. 2024a. Two-Step SPLADE: Simple, Efficient and Effective Approximation of SPLADE. In Advances in Information Retrieval: 46th European Conference on Information Retrieval, ECIR 2024, Glasgow, UK, March 24–28, 2024, Proceedings, Part II (Glasgow, United Kingdom). Springer-Verlag, Berlin, Heidelberg, 349–363. https://doi.org/10.1007/978-3-031-56060-6_23
- Lassance et al. (2024b) Carlos Lassance, Hervé Déjean, Thibault Formal, and Stéphane Clinchant. 2024b. SPLADE-v3: New baselines for SPLADE.
- Lassance et al. (2023) Carlos Lassance, Simon Lupart, Hervé Déjean, Stéphane Clinchant, and Nicola Tonellotto. 2023. A Static Pruning Study on Sparse Neural Retrievers. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval (Taipei, Taiwan) (SIGIR ’23). Association for Computing Machinery, New York, NY, USA, 1771–1775. https://doi.org/10.1145/3539618.3591941
@@
- Lin et al. (2021b) Sheng-Chieh Lin, Jheng-Hong Yang, and Jimmy Lin. 2021b. In-Batch Negatives for Knowledge Distillation with Tightly-Coupled Teachers for Dense Retrieval. In Proceedings of the 6th Workshop on Representation Learning for NLP (RepL4NLP-2021). Association for Computational Linguistics, Online, 163–173. https://doi.org/10.18653/v1/2021.repl4nlp-1.17
@@
- Lupart and Clinchant (2023) Simon Lupart and Stéphane Clinchant. 2023. A Study on FGSM Adversarial Training for Neural Retrieval. In Advances in Information Retrieval: 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2–6, 2023, Proceedings, Part II (Dublin, Ireland). Springer-Verlag, Berlin, Heidelberg, 484–492. https://doi.org/10.1007/978-3-031-28238-6_39
@@
- Mao et al. (2023a) Kelong Mao, Zhicheng Dou, Bang Liu, Hongjin Qian, Fengran Mo, Xiangli Wu, Xiaohua Cheng, and Zhao Cao. 2023a. Search-Oriented Conversational Query Editing. In Findings of the Association for Computational Linguistics: ACL 2023, Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 4160–4172. https://doi.org/10.18653/v1/2023.findings-acl.256
- Mao et al. (2023b) Kelong Mao, Zhicheng Dou, Fengran Mo, Jiewen Hou, Haonan Chen, and Hongjin Qian. 2023b. Large Language Models Know Your Contextual Search Intent: A Prompting Framework for Conversational Search. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 1211–1225. https://doi.org/10.18653/v1/2023.findings-emnlp.86
- Mao et al. (2023c) Kelong Mao, Hongjin Qian, Fengran Mo, Zhicheng Dou, Bang Liu, Xiaohua Cheng, and Zhao Cao. 2023c. Learning Denoised and Interpretable Session Representation for Conversational Search. In Proceedings of the ACM Web Conference 2023 (Austin, TX, USA) (WWW ’23). Association for Computing Machinery, New York, NY, USA, 3193–3202. https://doi.org/10.1145/3543507.3583265
@@
- Mo et al. (2023) Fengran Mo, Kelong Mao, Yutao Zhu, Yihong Wu, Kaiyu Huang, and Jian-Yun Nie. 2023. ConvGQR: Generative Query Reformulation for Conversational Search. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 4998–5012. https://doi.org/10.18653/v1/2023.acl-long.274
- Mo et al. (2024b) Fengran Mo, Chen Qu, Kelong Mao, Yihong Wu, Zhan Su, Kaiyu Huang, and Jian-Yun Nie. 2024b. Aligning Query Representation with Rewritten Query and Relevance Judgments in Conversational Search. arXiv preprint arXiv:2407.20189 (2024). arXiv:2407.20189 [cs.IR] https://arxiv.org/abs/2407.20189
@@
- Nogueira and Cho (2019) Rodrigo Nogueira and Kyunghyun Cho. 2019. Passage Re-ranking with BERT. arXiv preprint arXiv:1901.04085 (2019).
@@
- Qu et al. (2020) Chen Qu, Liu Yang, Cen Chen, Minghui Qiu, W. Bruce Croft, and Mohit Iyyer. 2020. Open-Retrieval Conversational Question Answering. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval (Virtual Event, China) (SIGIR ’20). Association for Computing Machinery, New York, NY, USA, 539–548. https://doi.org/10.1145/3397271.3401110
- Radlinski and Craswell (2017) Filip Radlinski and Nick Craswell. 2017. A Theoretical Framework for Conversational Search. In Proceedings of the 2017 Conference on Conference Human Information Interaction and Retrieval (Oslo, Norway) (CHIIR ’17). Association for Computing Machinery, New York, NY, USA, 117–126. https://doi.org/10.1145/3020165.3020183
@@
- Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych. 2019. Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics. http://arxiv.org/abs/1908.10084
- Robertson (2009) S. Robertson. 2009. The Probabilistic Relevance Framework: BM25 and Beyond. Foundations and Trends® in Information Retrieval 3, 4 (2009), 333–389. http://scholar.google.de/scholar.bib?q=info:U4l9kCVIssAJ:scholar.google.com/&output=citation&hl=de&as_sdt=2000&as_vis=1&ct=citation&cd=1
@@
- Tao and Zhai (2007) Tao Tao and ChengXiang Zhai. 2007. An exploration of proximity measures in information retrieval. In Proceedings of the 30th Annual International ACM SIGIR Conference on Research and Development in Information Retrieval (Amsterdam, The Netherlands) (SIGIR ’07). Association for Computing Machinery, New York, NY, USA, 295–302. https://doi.org/10.1145/1277741.1277794
@@
- Wolf et al. (2020) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Remi Louf, Morgan Funtowicz, Joe Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien Plu, Canwen Xu, Teven Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest, and Alexander Rush. 2020. Transformers: State-of-the-Art Natural Language Processing. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Qun Liu and David Schlangen (Eds.). Association for Computational Linguistics, Online, 38–45. https://doi.org/10.18653/v1/2020.emnlp-demos.6
@@
- Xiao et al. (2023) Yan Xiao, Yixing Fan, Ruqing Zhang, and Jiafeng Guo. 2023. Beyond Precision: A Study on Recall of Initial Retrieval with Neural Representations. In Information Retrieval: 28th China Conference, CCIR 2022, Chongqing, China, September 16–18, 2022, Revised Selected Papers (Chongqing, China). Springer-Verlag, Berlin, Heidelberg, 76–89. https://doi.org/10.1007/978-3-031-24755-2_7
@@
- Ye et al. (2023) Fanghua Ye, Meng Fang, Shenghui Li, and Emine Yilmaz. 2023. Enhancing Conversational Search: Large Language Model-Aided Informative Query Rewriting. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 5985–6006. https://doi.org/10.18653/v1/2023.findings-emnlp.398
- Yu et al. (2021) Shi Yu, Zhenghao Liu, Chenyan Xiong, Tao Feng, and Zhiyuan Liu. 2021. Few-Shot Conversational Dense Retrieval. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval (Virtual Event, Canada) (SIGIR ’21). Association for Computing Machinery, New York, NY, USA, 829–838. https://doi.org/10.1145/3404835.3462856
- Zamani et al. (2018) Hamed Zamani, Mostafa Dehghani, W. Bruce Croft, Erik Learned-Miller, and Jaap Kamps. 2018. From Neural Re-Ranking to Neural Ranking: Learning a Sparse Representation for Inverted Indexing. In Proceedings of the 27th ACM International Conference on Information and Knowledge Management (Torino, Italy) (CIKM ’18). Association for Computing Machinery, New York, NY, USA, 497–506. https://doi.org/10.1145/3269206.3271800
Appendix A Learned Sparse Retrieval
We recall in this section the basics of Information Retrieval and Learned Sparse Retrieval, with associated losses. InfoNCE (Oord et al., 2018) loss on a triplets , containing query, positive and negative documents:
We further recall the regularization FLOPS loss, defined from the mean probability within a batch for the token of the vocabulary to be activated in the representation.
Appendix B MSE-margin
Definition MSE-margin. While our distillation uses the Kullback Leibler Divergence, it is also possible to distill using the MSE-margin loss on the scores. The associated loss is given below:
The intuition is the same, but it emphasizes the margin between documents, rather than actual scores, as proposed by Hofstätter et al. (2020).
| Loss | QReCC | CAsT 2020 | ||
| R@100 | MRR | R@100 | MRR | |
| SPLADE Human | 0.912 | 0.448 | 0.646 | 0.636 |
| DiSCo KLD | 0.927 | 0.483 | 0.523 | 0.455 |
| DiSCo MSE-margin | 0.924 | 0.485 | 0.520 | 0.458 |
Experiments. In Table 4 we compare MSE-margin with KL Divergence. Both seem to have similar performance, showing the robustness of the training objective. One of the main benefits of the MSE-margin is that it normalizes the distribution of scores when distilling a teacher model with a different range of scores into a student. In our case, both student and teacher are the same model, only the input differs, so the range of scores is very similar, hence the good performance with the KLD.
Appendix C Prompts
Below we list a one-shot prompt for rewriting with the LLMs. We experimented both in zero-shot and one-shot but decided to use the one-shot to have better teacher models. The example is taken out of the QReCC dataset.

### v2

@@ Acknowledgements.
- Lassance et al. (2024a) Carlos Lassance, Hervé Dejean, Stéphane Clinchant, and Nicola Tonellotto. 2024a. Two-Step SPLADE: Simple, Efficient and Effective Approximation of SPLADE. In Advances in Information Retrieval: 46th European Conference on Information Retrieval, ECIR 2024, Glasgow, UK, March 24–28, 2024, Proceedings, Part II (Glasgow, United Kingdom). Springer-Verlag, Berlin, Heidelberg, 349–363. doi:10.1007/978-3-031-56060-6_23
- Lassance et al. (2024b) Carlos Lassance, Hervé Déjean, Thibault Formal, and Stéphane Clinchant. 2024b. SPLADE-v3: New baselines for SPLADE. arXiv preprint arXiv:2403.06789 (2024).
- Lassance et al. (2023) Carlos Lassance, Simon Lupart, Hervé Déjean, Stéphane Clinchant, and Nicola Tonellotto. 2023. A Static Pruning Study on Sparse Neural Retrievers. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval (Taipei, Taiwan) (SIGIR ’23). Association for Computing Machinery, New York, NY, USA, 1771–1775. doi:10.1145/3539618.3591941
@@
- Lin et al. (2021b) Sheng-Chieh Lin, Jheng-Hong Yang, and Jimmy Lin. 2021b. In-Batch Negatives for Knowledge Distillation with Tightly-Coupled Teachers for Dense Retrieval. In Proceedings of the 6th Workshop on Representation Learning for NLP (RepL4NLP-2021). Association for Computational Linguistics, Online, 163–173. doi:10.18653/v1/2021.repl4nlp-1.17
@@
- Lupart et al. (2024) Simon Lupart, Zahra Abbasiantaeb, and Mohammad Aliannejadi. 2024. IRLab@ iKAT24: Learned Sparse Retrieval with Multi-aspect LLM Query Generation for Conversational Search. arXiv preprint arXiv:2411.14739 (2024).
- Lupart and Clinchant (2023) Simon Lupart and Stéphane Clinchant. 2023. A Study on FGSM Adversarial Training for Neural Retrieval. In Advances in Information Retrieval: 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2–6, 2023, Proceedings, Part II (Dublin, Ireland). Springer-Verlag, Berlin, Heidelberg, 484–492. doi:10.1007/978-3-031-28238-6_39
@@
- Mao et al. (2023a) Kelong Mao, Zhicheng Dou, Bang Liu, Hongjin Qian, Fengran Mo, Xiangli Wu, Xiaohua Cheng, and Zhao Cao. 2023a. Search-Oriented Conversational Query Editing. In Findings of the Association for Computational Linguistics: ACL 2023, Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 4160–4172. doi:10.18653/v1/2023.findings-acl.256
- Mao et al. (2023b) Kelong Mao, Zhicheng Dou, Fengran Mo, Jiewen Hou, Haonan Chen, and Hongjin Qian. 2023b. Large Language Models Know Your Contextual Search Intent: A Prompting Framework for Conversational Search. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 1211–1225. doi:10.18653/v1/2023.findings-emnlp.86
- Mao et al. (2023c) Kelong Mao, Hongjin Qian, Fengran Mo, Zhicheng Dou, Bang Liu, Xiaohua Cheng, and Zhao Cao. 2023c. Learning Denoised and Interpretable Session Representation for Conversational Search. In Proceedings of the ACM Web Conference 2023 (Austin, TX, USA) (WWW ’23). Association for Computing Machinery, New York, NY, USA, 3193–3202. doi:10.1145/3543507.3583265
@@
- Mo et al. (2024b) Fengran Mo, Kelong Mao, Ziliang Zhao, Hongjin Qian, Haonan Chen, Yiruo Cheng, Xiaoxi Li, Yutao Zhu, Zhicheng Dou, and Jian-Yun Nie. 2024b. A survey of conversational search. arXiv preprint arXiv:2410.15576 (2024).
- Mo et al. (2023) Fengran Mo, Kelong Mao, Yutao Zhu, Yihong Wu, Kaiyu Huang, and Jian-Yun Nie. 2023. ConvGQR: Generative Query Reformulation for Conversational Search. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 4998–5012. doi:10.18653/v1/2023.acl-long.274
- Mo et al. (2024c) Fengran Mo, Chen Qu, Kelong Mao, Yihong Wu, Zhan Su, Kaiyu Huang, and Jian-Yun Nie. 2024c. Aligning Query Representation with Rewritten Query and Relevance Judgments in Conversational Search. arXiv preprint arXiv:2407.20189 (2024). arXiv:2407.20189 [cs.IR] https://arxiv.org/abs/2407.20189
@@
- Qu et al. (2020) Chen Qu, Liu Yang, Cen Chen, Minghui Qiu, W. Bruce Croft, and Mohit Iyyer. 2020. Open-Retrieval Conversational Question Answering. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval (Virtual Event, China) (SIGIR ’20). Association for Computing Machinery, New York, NY, USA, 539–548. doi:10.1145/3397271.3401110
- Radlinski and Craswell (2017) Filip Radlinski and Nick Craswell. 2017. A Theoretical Framework for Conversational Search. In Proceedings of the 2017 Conference on Conference Human Information Interaction and Retrieval (Oslo, Norway) (CHIIR ’17). Association for Computing Machinery, New York, NY, USA, 117–126. doi:10.1145/3020165.3020183
@@
- Tao and Zhai (2007) Tao Tao and ChengXiang Zhai. 2007. An exploration of proximity measures in information retrieval. In Proceedings of the 30th Annual International ACM SIGIR Conference on Research and Development in Information Retrieval (Amsterdam, The Netherlands) (SIGIR ’07). Association for Computing Machinery, New York, NY, USA, 295–302. doi:10.1145/1277741.1277794
@@
- Wolf et al. (2020) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Remi Louf, Morgan Funtowicz, Joe Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien Plu, Canwen Xu, Teven Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest, and Alexander Rush. 2020. Transformers: State-of-the-Art Natural Language Processing. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Qun Liu and David Schlangen (Eds.). Association for Computational Linguistics, Online, 38–45. doi:10.18653/v1/2020.emnlp-demos.6
@@
- Xiao et al. (2023) Yan Xiao, Yixing Fan, Ruqing Zhang, and Jiafeng Guo. 2023. Beyond Precision: A Study on Recall of Initial Retrieval with Neural Representations. In Information Retrieval: 28th China Conference, CCIR 2022, Chongqing, China, September 16–18, 2022, Revised Selected Papers (Chongqing, China). Springer-Verlag, Berlin, Heidelberg, 76–89. doi:10.1007/978-3-031-24755-2_7
@@
- Ye et al. (2023) Fanghua Ye, Meng Fang, Shenghui Li, and Emine Yilmaz. 2023. Enhancing Conversational Search: Large Language Model-Aided Informative Query Rewriting. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 5985–6006. doi:10.18653/v1/2023.findings-emnlp.398
- Yu et al. (2021) Shi Yu, Zhenghao Liu, Chenyan Xiong, Tao Feng, and Zhiyuan Liu. 2021. Few-Shot Conversational Dense Retrieval. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval (Virtual Event, Canada) (SIGIR ’21). Association for Computing Machinery, New York, NY, USA, 829–838. doi:10.1145/3404835.3462856
- Zamani et al. (2018) Hamed Zamani, Mostafa Dehghani, W. Bruce Croft, Erik Learned-Miller, and Jaap Kamps. 2018. From Neural Re-Ranking to Neural Ranking: Learning a Sparse Representation for Inverted Indexing. In Proceedings of the 27th ACM International Conference on Information and Knowledge Management (Torino, Italy) (CIKM ’18). Association for Computing Machinery, New York, NY, USA, 497–506. doi:10.1145/3269206.3271800